OpenStudy (anonymous):

HELP PLEASE: If three times Linda's age is decreased by 36, the result is twice Linda's age. How old is Linda???

OpenStudy (heisenberg):

just set up a simple algebra equation. you have one unknown: linda's age. we'll call this x. \[3*x - 36 = 2*x\] do you know how to solve from here?

OpenStudy (anonymous):

That's exactly what I got and Im having a hard time trying to solve this!

OpenStudy (heisenberg):

so the trick here is to get all your variables on one side. get the x's on the left and everything else on the right. then combine like terms: 3*x - 2*x = 36 1*x = 36 your answer!

OpenStudy (anonymous):

3x-6=2x 3x-2x=6 x(3-2)=6 x=6/(3-2)=6

OpenStudy (anonymous):

you need to isolate the variable which is x

OpenStudy (anonymous):

WHoa, sorry abouit that-went too fast. Should be x(3-2)=36 which of course is x=36

OpenStudy (anonymous):

@jcntrl: I got 6 as well, and its not working out after I check it, unless Im checking it wrong!

OpenStudy (anonymous):

It's a typo on my part. I typed 6 when I'd meant to type 36.

OpenStudy (anonymous):

So the steps should be: 3x-36=2x 3x-2x=36 x(3-2)=36 x=6/(3-2) x=6

OpenStudy (anonymous):

Oh, ok! Thanks guys! I think I got it! I have to show work too! But i understand now! Damn, Linda's old! LOL!

OpenStudy (anonymous):

Thanks everyone!

OpenStudy (anonymous):

Don't just copy what I've shown here. Understand it. WOork through it backwards. Otherwise on exam day, you'll be hosed. Cheers.

OpenStudy (anonymous):

36 isn't very old :)

OpenStudy (anonymous):

I know! LOL! Just a corney joke! Ya know what, that's the exact answer I got before, but when i went to check it, it didnt come out right! Could you help me check the answer, please?

OpenStudy (anonymous):

What is it that you don't understand? To check the answer, I would go back to the original problem statement and see if my solution holds true. So let's test if my solution of L=36 is true. If three times Linda's age is decreased by 36, the result is twice Linda's age. How old is Linda??? Then is this statement true: 3L-36 = 2L 3*36-36 = 2*36.............substitute L=36 36(3-1) = 2*36...............factor 36 out 3-1 = 2...........................divide both sides by 36 2=2.................................identity. Therefore, the solution is valid.

OpenStudy (anonymous):

Sorry, I magically switched from using x as my variable to using L as my variable. It really doesn't matter--a duck by any other name is still a duck--as long as you can keep it all straight on paper or in your head. :)